# Inner crust of neutron stars: Unified description of structure and superconductivity

### v1

@@ B.2 Variational equations
where $y$ is either of the variables from the set $\{n,\,x,\,n_{no},\,r_{N},\,u,\,n_{sk},\,n_{dc}\}$. In order to reveal the physical significance of the variational equations we shall discuss in detail each of the equations.
@@ B.3 Optimum size of nucleus
The derivative
$\left.\frac{\partial(w_{\rm tot})}{\partial r_{N}}\right|_{n,x,n_{no},u,n_{sk},n_{dc}}=0$ (77)
with respect to the nucleus size $r_{N}$ at fixed $\{n,\,x,\,n_{no},\,u,\,n_{sk},\,n_{dc}\}$ implies that the energy per baryon is minimized by nucleus size while the number densities are fixed. Application of the derivative to Eq. (64) yields the result known as the nuclear virial theorem with a correction due to the principal curvature of the thin surface of the dense (nuclear) phase of the inner crust matter:
@@
Notice that there are no contributions from $w_{\mathrm{sk}}$ and $w_{\mathrm{dc}}$. From Eq. (78) we find the equation that determines $r_{N}$:
@@
The solution to Eq. (79) can be readily found using the methods of Ferrari and Cardano resulting in
@@ B.4 Chemical equilibrium
The derivative
$\left.\frac{\partial(w_{\rm tot})}{\partial x}\right|_{n,n_{no},r_{N},u,n_{sk},n_{dc}}=0$ (85)
@@
$-\frac{1}{un}\left.\frac{\partial(w_{\mathrm{p.surf}}+w_{\mathrm{C+L}}+w_{\mathrm{curv}})}{\partial x}\right|_{n,n_{no},r_{N},u,n_{sk},n_{dc}}.$
Notice that there are no contributions from $w_{\mathrm{sk}}$ and $w_{\mathrm{dc}}$. Here, the electron chemical potential $\mu_{e}$ is:
$\mu_{e}=\frac{1}{un}\left.\frac{\partial w_{\mathrm{e}}}{\partial x}\right|_{n,n_{no},r_{N},u,n_{sk},n_{dc}}=\hbar c(3\pi^{2}uxn)^{\frac{1}{3}}.$ (87)
@@ B.5 Continuity of neutron chemical potential across the nucleus interface
The derivative
$\left.\frac{\partial(w_{\rm tot})}{\partial n_{ni}}\right|_{N_{n},n_{pi},r_{N},u,n_{sk},n_{dc}}=0$ (89)
@@
The partial derivative is given by:
$\left.\frac{\partial}{\partial n_{ni}}\right|_{N_{n},n_{pi},r_{N},u,n_{sk},n_{dc}}=$ (92)
$\left.\frac{\partial}{\partial n}\right|_{x,n_{no},r_{N},u,n_{sk},n_{dc}}+\frac{\delta n}{\delta x}\left.\frac{\partial}{\partial x}\right|_{n,n_{no},r_{N},u,n_{sk},n_{dc}}$
$+\frac{\delta n_{no}}{\delta n_{ni}}\left.\frac{\partial}{\partial n_{no}}\right|_{n,x,r_{N},u,n_{sk},n_{dc}}.$
@@
$\left.\frac{\partial}{\partial n_{ni}}\right|_{N_{n},n_{pi},r_{N},u,n_{sk},n_{dc}}=$ (93)
$\left.\frac{\partial}{\partial n}\right|_{x,\,n_{no},r_{N},u,n_{sk},n_{dc}}-\frac{x}{n}\left.\frac{\partial}{\partial x}\right|_{n,n_{no},r_{N},u,n_{sk},n_{dc}}$
$-\frac{u}{1-u}\left.\frac{\partial}{\partial n_{no}}\right|_{n,x,r_{N},u,n_{sk},n_{dc}}.$
@@
$-\frac{x}{un}\left.\frac{\partial(w_{\rm p.surf}+w_{\rm curv})}{\partial x}\right|_{r_{N},u,n_{sk},n_{dc}},$ (95)
@@
Notice that Eqs. (95) and (96) do not contain the contribution from $w_{\rm C+L}$ by virtue of Eq. (90). Moreover, there are no contributions from $w_{\mathrm{sk}}$ and $w_{\mathrm{dc}}$. The derivatives within $\mu_{ni}$ are calculated as following:
@@
$+\frac{1-x}{un}\left.\frac{\partial(w_{\rm p.surf}+w_{\rm curv})}{\partial x}\right|_{r_{N},u,n_{sk},n_{dc}}$
$+\frac{1}{un}\left.\frac{\partial(w_{\rm C+L})}{\partial x}\right|_{n,r_{N},u,n_{sk},n_{dc}}.$ (99)
@@ B.6 Continuity of pressure across the nucleus interface
The derivative
$\left.\frac{\partial(w_{\rm tot})}{\partial u}\right|_{N_{n},N_{no},x,V_{c},n_{sk},n_{dc}}=0$ (101)
implies that the energy density is minimized by choosing the optimal size of the nucleus while keeping fixed the cell size, the number of neutrons inside the nucleus, the number of neutrons outside the nucleus and the proton fraction. As a result, the pressure in the nucleus is equal to the pressure in the dripped neutrons, $P_{i}=P_{o}$.
@@
The partial derivative is given by:
$\left.\frac{\partial}{\partial u}\right|_{N_{n},N_{no},x,V_{c},n_{sk},n_{dc}}=$ (104)
$\left.\frac{\partial}{\partial u}\right|_{n,x,n_{no},V_{c},n_{sk},n_{dc}}+\frac{\delta n}{\delta u}\left.\frac{\partial}{\partial n}\right|_{n_{no},x,r_{N},u,n_{sk},n_{dc}}$
$+\frac{\delta n_{no}}{\delta u}\left.\frac{\partial}{\partial n_{no}}\right|_{n,x,r_{N},u,n_{sk},n_{dc}}.$
@@
$\left.\frac{\partial}{\partial u}\right|_{N_{n},N_{no},x,V_{c},n_{sk},n_{dc}}=$ (105)
$\left.\frac{\partial}{\partial u}\right|_{n,x,n_{no},V_{c},n_{sk},n_{dc}}-\frac{n}{u}\left.\frac{\partial}{\partial n}\right|_{n_{no},x,r_{N},u,n_{sk},n_{dc}}$
$+\frac{n_{no}}{1-u}\left.\frac{\partial}{\partial n_{no}}\right|_{n,x,r_{N},u,n_{sk},n_{dc}}.$
@@
$P_{i}=n^{2}\left.\frac{\partial\varepsilon}{\partial n}\right|_{x}-\left(\left.\frac{\partial}{\partial u}\right|_{n,x,n_{no},V_{c},n_{sk},n_{dc}}-\frac{n}{u}\left.\frac{\partial}{\partial n}\right|_{n_{no},r_{N},u,n_{sk},n_{dc}}\right)$
×$\times(w_{\rm p.surf}+w_{\rm C+L}+w_{\rm curv}),$ (107)
@@
$P_{o}=\frac{n_{no}}{1-u}\left.\frac{\partial w_{\mathrm{no}}}{\partial n_{no}}\right|_{u,n_{sk},n_{dc}}-\frac{w_{\mathrm{no}}}{1-u}\\ =n_{no}^{2}\frac{\partial\varepsilon(n_{no},0)}{\partial n_{no}}.$ (108)
Notice that there are no contributions from $w_{\mathrm{sk}}$ and $w_{\mathrm{dc}}$. With the help of Eq. (102) and the relations
@@
$\left.\frac{\partial w_{\rm p.surf}}{\partial u}\right|_{N_{n},N_{no},x,V_{c},n_{sk},n_{dc}}=\frac{d-1}{r_{N}}\sigma_{s}(x),$ (112)
$\left.\frac{\partial w_{\rm curv}}{\partial u}\right|_{N_{n},N_{no},x,V_{c},n_{sk},n_{dc}}=\frac{(d-2)(d-1)}{r_{N}^{2}}\sigma_{c}(x),$ (113)
@@
$\left(\left.\frac{\partial}{\partial u}\right|_{n,x,n_{no},V_{c},n_{sk},n_{dc}}-\frac{n}{u}\left.\frac{\partial}{\partial n}\right|_{n_{no},r_{N},u,n_{sk},n_{dc}}\right)w_{\rm C+L}$
@@
B.7 Optimal energy to add neutron to the skin or to the density curvature region
The condition
$\left.\frac{\partial(E_{\rm tot}/V_{c})}{\partial n_{sk}}\right|_{n,x,N_{n},r_{N},u,n_{dc}}=0$ (115)
implies that the energy density is minimized by the optimal value of the neutrons in the skin while keeping $(N_{sk}+N_{no})$ fixed. This condition induces the following relation:
$\delta n_{sk}+(1-u)\delta n_{no}=0.$ (116)
The partial derivative is given by
$\left.\frac{\partial}{\partial n_{sk}}\right|_{n,x,N_{n},r_{N},u,n_{dc}}=$ (117)
$\left.\frac{\partial}{\partial n_{sk}}\right|_{n,x,n_{no},V_{c},n_{dc}}+\left.\frac{\partial}{\partial n_{no}}\right|_{n,x,V_{c},n_{sk},n_{dc}}.$
Analogously, the condition
$\left.\frac{\partial(E_{\rm tot}/V_{c})}{\partial n_{dc}}\right|_{n,x,N_{n},r_{N},u,n_{sk}}=0$ (118)
implies that the energy density is minimized by the optimal value of the neutrons in the density curvature region while keeping $(N_{dc}+N_{no})$ fixed. This condition induces the following relation:
$\delta n_{dc}+(1-u)\delta n_{no}=0.$ (119)
The partial derivative is given by
$\left.\frac{\partial}{\partial n_{dc}}\right|_{n,x,N_{n},r_{N},u,n_{sk}}=$ (120)
$\left.\frac{\partial}{\partial n_{dc}}\right|_{n,x,n_{no},V_{c},n_{sk}}+\left.\frac{\partial}{\partial n_{no}}\right|_{n,x,V_{c},n_{sk},n_{dc}}.$
Application of these partial derivatives to Eqs. (187) and (188) yields
$\mu_{sk}=\mu_{no},$ (122)
$\mu_{dc}=\mu_{no}.$
Appendix C Nuclear bubbles in LDM
The independent variables of LDM for the bubble phases are
$\{n_{b},\,n,\,x,\,r_{B},\,n_{no},\,n_{sk}^{\rm bub},\,n_{dc}^{\rm bub}\},$ (123)
where the baryon density and the proton fraction outside the bubble $n$ and $x$, correspondingly, the bubble radius $r_{B}$, the neutron skin density $n_{sk}^{\rm bub}$ and the curvature density $n_{dc}^{\rm bub}$ are defined below. As in the case with nuclear clusters, in the practical problem the unknown variables are:
$\{n,\,x,\,r_{B},\,n_{no},\,n_{sk}^{\rm bub},\,n_{dc}^{\rm bub}\},$ (124)
@@ C.1 Energy density
where $r_{c}$ is the radius of the unit cell and $d$ is the dimensionality ($d=3$ for 3B, $d=2$ for 2B and $d=1$ for 1B). The volume of the unit cell is
$V_{c}=\left\{\begin{array}[]{c}({4}/{3})\pi r_{c}^{3},\quad(3B)\\ \pi r_{c}^{2}\times L,\quad(2B)\\ 2r_{c}\times L^{2},\quad(1B)\end{array}\right.$ (128)
@@
which corresponds to neutrons outside the bubble, at the surface and inside the bubble. We define the following neutron number densities.
@@
The surface neutrons (adsorbed on the surface of the nucleus) can be represented as neutrons associated with the neutron skin and with the surface width (the density curvature):
$N_{surf}=N_{sk}+N_{dc}.$ (134)
The cell-averaged number density of the skin neutrons is
$n_{sk}^{\rm bub}=\frac{N_{sk}}{V_{c}}.$ (135)
The cell-averaged number density of neutrons associated with the density curvature is
$n_{dc}^{\rm bub}=\frac{N_{dc}}{V_{c}}.$ (136)
Derivation of the governing variational equations of LDM assumes that $n_{sk}^{\rm bub}$ and $n_{dc}^{\rm bub}$ are independent variables. It is important that as soon as the variational equations are derived, the quantities $n_{sk}^{\rm bub}$ and $n_{dc}^{\rm bub}$ can be specified from the microscopic equations of the Skyrme model as following:
$n_{sk}^{\rm bub}=\left\{\begin{array}[]{c}{4\pi r_{B}^{2}s_{n}(n_{L}-n_{G})}{V_{c}^{-1}},\quad(3B)\\ {2\pi r_{B}s_{n}(n_{L}-n_{G})}{V_{c}^{-1}}\times L,\quad(2B)\\ {s_{n}(n_{L}-n_{G})}{V_{c}^{-1}}\times L^{2},\quad(1B)\end{array}\right.$ (137)
where $s_{n}=s_{n}(x)$ is the neutron skin thickness, and
$n_{dc}^{\rm bub}=\left\{\begin{array}[]{c}{8\pi r_{B}(b_{n}^{2}-b_{p}^{2}+s_{n}^{2})(n_{L}-n_{G})}{V_{c}^{-1}},\quad(3B)\\ {2\pi(b_{n}^{2}-b_{p}^{2}+s_{n}^{2})(n_{L}-n_{G})}{V_{c}^{-1}}\times L,\quad(2B)\\ 0,\quad(1N)\end{array}\right.$ (138)
where $b_{n}=b_{n}(x)$, $b_{p}=b_{p}(x)$ are the neutron surface width, proton surface width and $n_{L}=n_{L}(x)$ and $n_{G}=n_{G}(x)$ are some parameters, which can be found from the Skyrme model and will be presented elsewhere.
@@
In our model, the total energy density includes 8 terms:
$w_{\mathrm{tot}}^{\rm bub}=w_{\mathrm{nuc}}^{\rm bub}+w_{\mathrm{p.surf}}^{\rm bub}+w_{\mathrm{C+L}}^{\rm bub}+w_{\mathrm{no}}^{\rm bub}+w_{\mathrm{e}}^{\rm bub}+w_{\mathrm{curv}}^{\rm bub}+w_{\mathrm{sk}}^{\rm bub}+w_{\mathrm{dc}}^{\rm bub}.$ (141)
The rest mass contribution due to baryons and the strong interaction contribution associated with the uniform matter inside the nucleus is given by
$w_{\rm nuc}=(1-u^{\rm bub})n\left[\left(1-x\right)m_{n}+xm_{p}\right]c^{2}+(1-u^{\rm bub})n\varepsilon(n,x),$ (142)
where $\varepsilon$ is the energy per baryon specified in the Appendix.
@@
where the function $f_{d}(u)$ is given by Eq. (69).
@@
The energy density associated with curvature of the nucleus surface $w_{\mathrm{curv}}^{\rm bub}$ is proportional to both the surface area of the bubble and to the mean principal curvature of the surface of the nucleus, which is $2/r_{B}$ for 3D clusters, $1/r_{B}$ for 2D clusters and 0 for 1D clusters. Thus, it has the form
@@
The energy density associated with the neutron skin can be written as
$w_{\mathrm{sk}}^{\rm bub}=n_{\mathrm{sk}}^{\rm bub}\mu_{\mathrm{sk}},$ (148)
where $\mu_{\mathrm{sk}}$ is defined by the variational equation with respect to $n_{\mathrm{sk}}^{\rm bub}$.
Analogously, the energy density associated with the nuclear density curvature can be written as
$w_{\mathrm{dc}}^{\rm bub}=n_{\mathrm{dc}}^{\rm bub}\mu_{\mathrm{dc}},$ (149)
where $\mu_{\mathrm{dc}}$ is defined by the variational equation with respect to $n_{\mathrm{dc}}^{\rm bub}$.

### v2

@@ B.2 Variational equations
where $y$ is either of the variables from the set $\{n,\,x,\,n_{no},\,r_{N},\,u\}$. In order to reveal the physical significance of the variational equations we shall discuss in detail each of the equations.
@@ B.3 Optimum size of nucleus
The extremum
$\left.\frac{\partial(w_{\rm tot})}{\partial r_{N}}\right|_{n,x,n_{no},u}=0$ (81)
with respect to the nucleus size $r_{N}$ at fixed $\{n,\,x,\,n_{no},\,u\}$ implies that the energy per baryon is minimized by nucleus size while the number densities are fixed. Application of the derivative to Eq. (70) yields the result known as the nuclear virial theorem with a correction due to the principal curvature of the thin surface of the dense (nuclear) phase of the inner crust matter:
@@
Notice that in Eq. (82) there are no contributions from neutron skin even if the corresponding energy density were included into $w_{\rm{tot}}$. From Eq. (82) we find the equation that determines $r_{N}$:
@@
The solution to Eq. (83) can be readily found using the well-known method Cardano1545 resulting in
@@ B.4 Chemical equilibrium
The extremum
$\left.\frac{\partial(w_{\rm tot})}{\partial x}\right|_{n,n_{no},r_{N},u}=0$ (89)
@@
$-\frac{1}{un}\left.\frac{\partial(w_{\mathrm{p.surf}}+w_{\mathrm{C+L}}+w_{\mathrm{curv}})}{\partial x}\right|_{n,n_{no},r_{N},u}.$
Notice that, again, there are no contributions from neutron skin even if the corresponding energy density were included into $w_{\rm{tot}}$. Here, the electron chemical potential $\mu_{e}$ is:
$\mu_{e}=\frac{1}{un}\left.\frac{\partial w_{\mathrm{e}}}{\partial x}\right|_{n,n_{no},r_{N},u}=\hbar c(3\pi^{2}uxn)^{\frac{1}{3}}.$ (91)
@@ B.5 Continuity of neutron chemical potential across the nucleus interface
The extremum
$\left.\frac{\partial(w_{\rm tot})}{\partial n_{ni}}\right|_{N_{n},n_{pi},r_{N},u}=0$ (93)
@@
The partial derivative reads
$\left.\frac{\partial}{\partial n_{ni}}\right|_{N_{n},n_{pi},r_{N},u}=$ (96)
$\left.\frac{\partial}{\partial n}\right|_{x,n_{no},r_{N},u}+\frac{\delta n}{\delta x}\left.\frac{\partial}{\partial x}\right|_{n,n_{no},r_{N},u}$
$+\frac{\delta n_{no}}{\delta n_{ni}}\left.\frac{\partial}{\partial n_{no}}\right|_{n,x,r_{N},u}.$
@@
$\left.\frac{\partial}{\partial n_{ni}}\right|_{N_{n},n_{pi},r_{N},u}=$ (97)
$\left.\frac{\partial}{\partial n}\right|_{x,\,n_{no},r_{N},u}-\frac{x}{n}\left.\frac{\partial}{\partial x}\right|_{n,n_{no},r_{N},u}$
$-\frac{u}{1-u}\left.\frac{\partial}{\partial n_{no}}\right|_{n,x,r_{N},u}.$
@@
$-\frac{x}{un}\left.\frac{\partial(w_{\rm p.surf}+w_{\rm curv})}{\partial x}\right|_{r_{N},u},$ (99)
@@
Notice that Eqs. (99) and (100) do not contain the contribution from $w_{\rm C+L}$ by virtue of Eq. (94). Moreover, there are no contributions from neutron skin even if the corresponding energy density were included into $w_{\rm{tot}}$. The derivatives within $\mu_{ni}$ are calculated as following:
@@
$+\frac{1-x}{un}\left.\frac{\partial(w_{\rm p.surf}+w_{\rm curv})}{\partial x}\right|_{r_{N},u}$
$+\frac{1}{un}\left.\frac{\partial(w_{\rm C+L})}{\partial x}\right|_{n,r_{N},u}.$ (103)
@@ B.6 Continuity of pressure across the nucleus interface
The extremum
$\left.\frac{\partial(w_{\rm tot})}{\partial u}\right|_{N_{n},N_{no},x,V_{c}}=0$ (105)
implies that the energy density is minimized by choosing the optimal size of the nucleus while keeping fixed the cell size, the number of neutrons, both inside and outside the nucleus, and the proton fraction. As a result, the pressure in the nucleus is equal to the pressure in the dripped neutrons, $P_{i}=P_{o}$.
@@
The partial derivative is reads
$\left.\frac{\partial}{\partial u}\right|_{N_{n},N_{no},x,V_{c}}=$ (108)
$\left.\frac{\partial}{\partial u}\right|_{n,x,n_{no},V_{c}}+\frac{\delta n}{\delta u}\left.\frac{\partial}{\partial n}\right|_{n_{no},x,r_{N},u}$
$+\frac{\delta n_{no}}{\delta u}\left.\frac{\partial}{\partial n_{no}}\right|_{n,x,r_{N},u}.$
@@
$\left.\frac{\partial}{\partial u}\right|_{N_{n},N_{no},x,V_{c}}=$ (109)
$\left.\frac{\partial}{\partial u}\right|_{n,x,n_{no},V_{c}}-\frac{n}{u}\left.\frac{\partial}{\partial n}\right|_{n_{no},x,r_{N},u}$
$+\frac{n_{no}}{1-u}\left.\frac{\partial}{\partial n_{no}}\right|_{n,x,r_{N},u}.$
@@
$P_{i}=n^{2}\left.\frac{\partial\varepsilon}{\partial n}\right|_{x}-\left(\left.\frac{\partial}{\partial u}\right|_{n,x,n_{no},V_{c}}-\frac{n}{u}\left.\frac{\partial}{\partial n}\right|_{n_{no},r_{N},u}\right)$
$\times(w_{\rm p.surf}+w_{\rm C+L}+w_{\rm curv}),$ (111)
@@
$P_{o}=\frac{n_{no}}{1-u}\left.\frac{\partial w_{\mathrm{no}}}{\partial n_{no}}\right|_{u}-\frac{w_{\mathrm{no}}}{1-u}\\ =n_{no}^{2}\frac{\partial\varepsilon(n_{no},0)}{\partial n_{no}}.$ (112)
Notice that there are no contributions from neutron skin even if the corresponding energy density were included into $w_{\rm{tot}}$. With the help of Eq. (106) and the relations
@@
$\left.\frac{\partial w_{\rm p.surf}}{\partial u}\right|_{N_{n},N_{no},x,V_{c}}=\frac{d-1}{r_{N}}\sigma_{s}(x),$ (116)
$\left.\frac{\partial w_{\rm curv}}{\partial u}\right|_{N_{n},N_{no},x,V_{c}}=\frac{(d-2)(d-1)}{r_{N}^{2}}\sigma_{c}(x),$ (117)
@@
$\left(\left.\frac{\partial}{\partial u}\right|_{n,x,n_{no},V_{c}}-\frac{n}{u}\left.\frac{\partial}{\partial n}\right|_{n_{no},r_{N},u}\right)w_{\rm C+L}$
@@
Appendix C Bubble phases in liquid drop model
A complete set of independent variables for the bubble phases 1B, 2B, 3B are
$\{n_{b},\,n,\,x,\,r_{B},\,n_{no}\},$ (119)
where the baryon density and the proton fraction outside the bubble $n$ and $x$, correspondingly, the bubble radius $r_{B}$. As in the case with nuclear clusters, in the practical problem the unknown variables are:
$\{n,\,x,\,r_{B},\,n_{no}\},$ (120)
@@ C.1 Energy density
where $r_{c}$ is the radius of the unit cell, $d=3$ for 3B and $d=2$ for 2B.
@@
which corresponds to neutrons outside the bubble, at the surface and inside the bubble. Again, we assume $N_{surf}=0$. We define the following neutron number densities.
@@
In our model, the total energy density includes 6 terms:
$w_{\mathrm{tot}}^{\rm bub}=w_{\mathrm{nuc}}^{\rm bub}+w_{\mathrm{p.surf}}^{\rm bub}+w_{\mathrm{C+L}}^{\rm bub}+w_{\mathrm{no}}^{\rm bub}+w_{\mathrm{e}}^{\rm bub}+w_{\mathrm{curv}}^{\rm bub}.$ (131)
The rest mass contribution due to baryons and the strong interaction contribution associated with the uniform matter outside the bubble reads
$w_{\rm nuc}^{\rm bub}=(1-u^{\rm bub})n\left[\left(1-x\right)m_{n}+xm_{p}\right]c^{2}+(1-u^{\rm bub})n\varepsilon(n,x),$ (132)
where $\varepsilon$ is the energy per baryon specified in Eq. (47).
@@
where the function $f_{d}(u)$ is given in Eq. (75).
@@
The mean principal curvature of the bubble surface is $2/r_{B}$ for bubbles of 3B and $1/r_{B}$ for bubbles of 2B. Thus, the energy density associated with curvature of the nucleus surface has the form